# Paul-Erik Haacker, Remco I. Leine, Renu Chaudhary, Kai Diethelm, André Schmidt, Safoura Hashemishahraki

### v2

@@ 1 Introduction
Recently, interest in fractional-order differential equations has grown in various fields. A number of dynamical processes with memory have proven to be better described using fractional-order differential equations, where integer-order models fail to provide low-degree-of-freedom representations. Among others, applications include viscoelastic materials, diffusion processes and control theory [Das2011]. Such – possibly nonlinear – systems admit periodic solutions when periodically forced or in the presence of self-excitation. Common formulations of Caputo differential equations cannot have exactly periodic solutions [Tavazoei2009], which is closely related to the fact that Caputo derivatives and Riemann-Liouville integrals of periodic functions are only asymptotically periodic [Area2016]. Studies [Rand2010, Daibiri2015] concerned with finding steady-state solutions of differential equations involving Liouville-Weyl-type fractional-order derivatives have successfully applied harmonic balance methods. A linearization of the system dynamics about a periodic solution, describing the dynamics of small perturbations, yields a LTP-FODE. This motivates the study of the stability of such systems.
@@
Lyapunov-stability of equilibria and $\Omega$-limit sets is a central concept in the study of dynamical systems and has also been studied for fractional order systems. Classically, stability of linear time-invariant fractional-order differential equations can be studied through Mittag-Leffler functions, see [Diethelm2010] and the references cited therein. As fractional-order systems typically admit algebraic decay [CongEtAl2020], the notion of Mittag-Leffler stability was introduced analogously to exponential stability [LiEtAl2009]. Furthermore, as fractional-order derivatives are nonlocal, FODEs are similar to functional differential equations [Hale1977]. This similarity has been exploited by [Hinze2020, Hinze2020a] in mechanical systems with fractional-order damping, treating the system as a functional differential equation with infinite delay and finding Lyapunov-functionals through the infinite-state representation (also called diffusive representation or kernel compression), which is often used for efficient numerical approximations of fractional-order integrals [Chaudhary2024]. LTI-FODEs perturbed by small nonlinearities were studied in [CongEtAl2018]. That work also analyzes linear time-varying FODEs, but only under the assumption of sufficiently small time-dependent terms. To the best of our knowledge, no stronger results on stability of solutions of linear time-variant fractional-order Caputo differential equations are available.
@@
In classical integer-order models (i.e. first-order ordinary differential equations), stability of periodic solutions may be analyzed via the eigenvalues of the infinite-dimensional Hill matrix which correspond to the Floquet exponents, see [Bayer2023a] and the references cited therein. These problems are understood quite deeply, with Floquet theory as their foundation. This theory may be applied in bifurcation analysis of periodic motion of mechanical systems, which is self-excited [Horvath2022] or externally forced [Youssef2025]. For LTP-FODEs, no similar theory has been established, although several attempts for the development of a Floquet theory for fractional-order systems have been undertaken [Rezazadeh2016, Iomin2023]. In particular, the work [Rezazadeh2016] assumes that the Mittag-Leffler function approximately satisfies a multiplicative identity, despite the fact that the associated error may become arbitrarily large. Further, in [Rezazadeh2016] a theory entirely based on approximations is developed, which only holds exactly in the first-order case ($\alpha=1$). However, parameter identification schemes applied to experimental data typically yield fractional orders between $0.1$ and $0.7$ [Pawlak2021, HinzeExperiment, Kim2009, Stark2021]. The work [Iomin2023] considers a Schrödinger LTP-FODE of Caputo-type. Therein, a solution form expressed by a series of Mittag-Leffler functions is considered; however, an analysis of asymptotic properties of the solution form as well as a Hill-type method to access characteristic multipliers is missing. Furthermore, the proposed form of solutions of [Iomin2023] uses a basis of Mittag–Leffler functions whereas the time-periodic system matrix (the Hamiltonian in the terminology of [Iomin2023]) is represented in harmonic form. The use of two different series impedes a Hill-type approach, since the Mittag-Leffler function and the exponential function do not share a multiplicative identity.
@@
To reach a Floquet theory, results on fundamental solution operators for LTP systems are lacking. The more general case of linear time-varying systems was considered by [Gomoyunov2020], developing a fundamental solution matrix. Defining a monodromy operator and analyzing its spectrum, as done for functional differential equations [Hale1977], remains an open challenge in the study of LTP-FODEs. Previous studies performed transition curve analysis on the fractional-order Mathieu equation [Rand2010, Shahroudi2024], as well as on more general forms  [Daibiri2015], also including first-order derivatives and time-delays. The latter work employs fractional-order derivatives with infinite memory, i.e. Liouville-Weyl-type derivatives, which is of particular relevance to our study. However, the analysis of transition curves in the parameter space only gives information on stability boundaries and thus fails to completely characterize stability over a parameter space, nor does it give information about the type of bifurcation which may originate at the stability boundary.
@@
The literature review shows, that the classical Caputo-type differential equations do not admit exactly periodic solutions, but do admit them as limit sets. In order to study the stability properties of such limit sets, one must first formulate an extension of the class of Caputo differential equations to also include periodic solutions. The Liouville-Weyl derivative has these properties and a rigorous formulation of an initial value problem along with the set of initial conditions has to be developed. This then opens paths to extend classical theories on periodic solutions to fractional-order systems. For example, fundamental tools such as the Poincaré–Bendixson Theorem [HirschEtAl2013], which provides conditions for the existence of periodic solutions, need to be developed in the fractional-order case. Further, when given a periodic solution, possibly obtained using the harmonic balance method [Rand2010, Daibiri2015, Shahroudi2024], linearly perturbed dynamics can be studied to assess stability in the hyperbolic case. Therefore, a Floquet-type theory on LTP-FODEs needs to be developed and the existence of Floquet exponents needs to be studied. For application of this theory to numerical stability analysis, methods like the Hill method should be extended to the fractional-order case.
@@
The main merit of this paper is the insight that exponentially decaying Floquet solutions (i.e. the ones related to asymptotically stable equilibria or stable manifolds of saddle equilibria) do not exist for LTP-FODEs, but exponentially growing Floquet solutions may be observed and assessed via a Hill-type approach. Parts of the research of this work were presented in a preliminary form in reference [Haacker2025], namely the fractional Hill problem was presented as well as results on linear time-invariant FODEs. The novel contributions of the present work are the formulation of the system class in Section 3, the analysis of decaying solutions in Section 4 as well as the proof of Theorem 5.1 and the following Propositions.
@@ 2.1 Fractional Calculus
We briefly recall some concepts of fractional calculus and refer the reader to the books [Diethelm2010, Kilbas2006] for further details. Several definitions of fractional derivatives exist, each suited to different mathematical and physical contexts. Given a function $f\in L^{1}[a,b]$, the Riemann–Liouville (RL) fractional integral of order $\alpha>0$ is defined as
@@
$\left({}^{\mathrm{RL}}\operatorname{D}_{a}^{\alpha}f\right)(t)=\left(\operatorname{D}^{1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}I_{a}^{1-\alpha}}f\right)(t).$ (2)
@@
To overcome this limitation of the RL derivative, the Caputo derivative is often used in applied settings [Hinze2020, Rand2010, Daibiri2015]. For $\alpha\in(0,1)$ and a function $f\in AC[a,b]$, it is defined as
@@
A different formulation (see [Diethelm2010] and the references therein) of the Caputo derivative reads
@@
for the case of $\alpha\in(0,1)$. If $f\in AC[a,b]$, then the two formulations coincide $\operatorname{D}_{\ast a}^{\alpha}f={}^{\mathrm{C}}\operatorname{D}_{a}^{\alpha}f$ almost everywhere [Diethelm2010].
@@ 2.2 The Hill Method for First-Order Systems
Hill’s method is a frequency-based approach to assess the stability of periodic solutions [Zhou2002]. It has been applied to first order ordinary differential equations as well as delay differential equations [Stepan2011]. Here, we summarize the approach for ordinary differential equations, which we develop for FODEs in Section 5. For time-periodic first-order systems
@@
Herein, $\tilde{J}_{k}$ are the Fourier coefficients of $\tilde{J}(t)=\sum_{k=-\infty}^{+\infty}\tilde{J}_{k}\exp(ik\omega t)$, $\omega=2\pi/T$ and $I$ is the identity matrix of appropriate dimension. After truncation to finite dimension $N$, one may either compute all eigenvalues of the truncated Hill-matrix $\mathbf{H}_{N}$ and employ a sorting technique to extract Floquet exponents or perform the recently introduced Koopman-Hill projection [Bayer2023a], avoiding sorting and directly approximating Floquet-multipliers.
@@ 3 Liouville-Weyl Differential Equations: A Class of Systems with Periodic Solutions
$X(-\infty,t_{0}]:=$ $\Bigl\{x_{0}\in{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}AC_{\mathrm{loc}}(-\infty,t_{0}]}\cap L^{\infty}(-\infty,t_{0}]~|~\exists\,\eta>0:x_{0}^{\prime}\in L^{\infty}[t_{0}-\eta,t_{0}]\Bigr\}$ (13)
@@
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{}^{\mathrm{C}}\operatorname{D}_{t_{0}}^{\alpha}x(t)$ $=\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}f(t,x(t))-\mathcal{F}x_{0}(t)=:f_{0}(t,x(t)),\quad t\geq t_{0},\quad x(t_{0})=x_{0}(t_{0}).$ (18)
Typically, the $\operatorname{D}_{\ast 0}^{\alpha}$-type differential equations are considered instead of $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}-$type formulations, since the latter shows problems on the space $C[0,h]$, but the two formulations coincide on $AC[0,h]$, see [Webb2025, Section 2.3]. We formalize this for our case in the Proposition 3.5 below. Consider the following example of $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}-$type problems on $C[0,h]$.
@@
Let $x:[0,1]\to\mathbb{R}$ be the Cantor ternary function [GO2003, pp. 96ff.]. This function is continuous, monotonically increasing, differentiable almost everywhere, and satisfies $x^{\prime}(t)=0$ a. e. Thus, $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}x(t)=0$ for $t\in[0,1]$ and $x$ is a strong solution of the initial value problem $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}x(t)=0$, $x(0)=0$. By a strong solution we mean a continuous function satisfying the differential equation and initial condition.
@@
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\operatorname{D}_{\ast t_{0}}^{\alpha}x(t)$ $=\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}f(t,x(t))-\mathcal{F}x_{0}(t)=:f_{0}(t,x(t)),\quad t\geq t_{0},\quad x(t_{0})=x_{0}(t_{0}).$ (19)
Since $f$ is continuous in its first variable and locally Lipschitz in its second variable, the same properties hold for $f_{0}$, because Lemma 3.3 guarantees that the forcing term $\mathcal{F}x_{0}$ is continuous. We thus have existence of strong continuous solutions in the sense of Theorem 6.5 of [Diethelm2010]. These solutions may or may not coincide with solutions of the $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}-$type problem (18). Of course, since the two operators $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}$ and $\operatorname{D}_{\ast 0}^{\alpha}$ coincide almost everywhere in the case of absolutely continuous functions, we have the following statement.
@@
Consider the Weierstrass function [Samko1994]
@@
which is continuous but nowhere differentiable on $[0,1]$. This function $W$ is Hölder continuous with exponent $\beta$ for all $\beta\in(0,1)$ and $W(0)=0$. Following the results of [Samko1994], $W$ thus has a continuous Caputo fractional derivative $\operatorname{D}_{\ast 0}^{\alpha}W$ of order $\alpha$ for all $\alpha\in(0,1)$. Then, the function $W$ can be seen as a solution of the initial-value problem
@@
The example (i) illustrates that exponentially decaying initial functions are not feasible, as these grow unboundedly in negative time. We will also observe in the next section that decaying solutions of linear versions of (12a) possess an algebraic decay instead of an exponential decay. Exponential growth in case (ii) is indeed possible and will be observed in the two following sections. In example (iii), the initialization function $x_{0}$ is a typical trajectory generated by a linear fractional-order Caputo differential equation with finite lower bound. Although its first derivative is undefined at time t=1t=-1, its forcing term is well defined. Example (iv) is an example showing that a singularity of $x_{0}^{\prime}$ at $t_{0}$ can lead to nonexistence of $\mathcal{F}x_{0}$. This particular example of $x_{0}$ is bounded and uniformly continuous, showing that the space of initial functions of (12a) needs to be chosen differently than in [Hinze2020]. Example (v) allows for an explicit computation of the forcing term. As this work is motivated by the study of periodic solutions, example (vi) is a periodic initial function. In fact, the Liouville-Weyl differential equation (12a) can possibly posess a periodic solution, in contrast to fractional differential equations with finite lower bound [Tavazoei2009]. Indeed, a fractional-order Caputo derivative with finite lower bound of a periodic function is not periodic [Area2016]. A minimal example of a system with a periodic solution is given in the following.
@@
We intend to study stability properties of linear cases of the system (12a) in the subsequent sections. We give the stability definition here, which is equivalent to the one of [Hinze2020] and the classical literature on functional differential equations [Hale1977].

### v3

@@ 1 Introduction
Recently, interest in fractional-order differential equations has grown in various fields. A number of dynamical processes with memory have proven to be better described using fractional-order differential equations, where integer-order models fail to provide low-degree-of-freedom representations. Among others, applications include viscoelastic materials, diffusion processes and control theory [7]. Such – possibly nonlinear – systems admit periodic solutions when periodically forced or in the presence of self-excitation. Common formulations of Caputo differential equations cannot have exactly periodic solutions [34], which is closely related to the fact that Caputo derivatives and Riemann-Liouville integrals of periodic functions are only asymptotically periodic [1]. Studies [28, 3] concerned with finding steady-state solutions of differential equations involving Liouville-Weyl-type fractional-order derivatives have successfully applied harmonic balance methods. A linearization of the system dynamics about a periodic solution, describing the dynamics of small perturbations, yields a LTP-FODE. This motivates the study of the stability of such systems.
@@
Lyapunov-stability of equilibria and $\Omega$-limit sets is a central concept in the study of dynamical systems and has also been studied for fractional order systems. Classically, stability of linear time-invariant fractional-order differential equations can be studied through Mittag-Leffler functions, see [10] and the references cited therein. As fractional-order systems typically admit algebraic decay [5], the notion of Mittag-Leffler stability was introduced analogously to exponential stability [26]. Furthermore, as fractional-order derivatives are nonlocal, FODEs are similar to functional differential equations [16]. This similarity has been exploited by [18, 17] in mechanical systems with fractional-order damping, treating the system as a functional differential equation with infinite delay and finding Lyapunov-functionals through the infinite-state representation (also called diffusive representation or kernel compression), which is often used for efficient numerical approximations of fractional-order integrals [4]. LTI-FODEs perturbed by small nonlinearities were studied in [6]. That work also analyzes linear time-varying FODEs, but only under the assumption of sufficiently small time-dependent terms. To the best of our knowledge, no stronger results on stability of solutions of linear time-variant fractional-order Caputo differential equations are available.
@@
In classical integer-order models (i.e. first-order ordinary differential equations), stability of periodic solutions may be analyzed via the eigenvalues of the infinite-dimensional Hill matrix which correspond to the Floquet exponents, see [2] and the references cited therein. These problems are understood quite deeply, with Floquet theory as their foundation. This theory may be applied in bifurcation analysis of periodic motion of mechanical systems, which is self-excited [21] or externally forced [37]. For LTP-FODEs, no similar theory has been established, although several attempts for the development of a Floquet theory for fractional-order systems have been undertaken [29, 23]. In particular, the work [29] assumes that the Mittag-Leffler function approximately satisfies a multiplicative identity, despite the fact that the associated error may become arbitrarily large. Further, in [29] a theory entirely based on approximations is developed, which only holds exactly in the first-order case ($\alpha=1$). However, parameter identification schemes applied to experimental data typically yield fractional orders between $0.1$ and $0.7$ [27, 19, 25, 33]. The work [23] considers a Schrödinger LTP-FODE of Caputo-type. Therein, a solution form expressed by a series of Mittag-Leffler functions is considered; however, an analysis of asymptotic properties of the solution form as well as a Hill-type method to access characteristic multipliers is missing. Furthermore, the proposed form of solutions of [23] uses a basis of Mittag–Leffler functions whereas the time-periodic system matrix (the Hamiltonian in the terminology of [23]) is represented in harmonic form. The use of two different series impedes a Hill-type approach, since the Mittag-Leffler function and the exponential function do not share a multiplicative identity.
@@
To reach a Floquet theory, results on fundamental solution operators for LTP systems are lacking. The more general case of linear time-varying systems was considered by [12], developing a fundamental solution matrix. Defining a monodromy operator and analyzing its spectrum, as done for functional differential equations [16], remains an open challenge in the study of LTP-FODEs. Previous studies performed transition curve analysis on the fractional-order Mathieu equation [28, 32], as well as on more general forms  [3], also including first-order derivatives and time-delays. The latter work employs fractional-order derivatives with infinite memory, i.e. Liouville-Weyl-type derivatives, which is of particular relevance to our study. However, the analysis of transition curves in the parameter space only gives information on stability boundaries and thus fails to completely characterize stability over a parameter space, nor does it give information about the type of bifurcation which may originate at the stability boundary.
@@
The literature review shows, that the classical Caputo-type differential equations do not admit exactly periodic solutions, but do admit them as limit sets. In order to study the stability properties of such limit sets, one must first formulate an extension of the class of Caputo differential equations to also include periodic solutions. The Liouville-Weyl derivative has these properties and a rigorous formulation of an initial value problem along with the set of initial conditions has to be developed. This then opens paths to extend classical theories on periodic solutions to fractional-order systems. For example, fundamental tools such as the Poincaré–Bendixson Theorem [20], which provides conditions for the existence of periodic solutions, need to be developed in the fractional-order case. Further, when given a periodic solution, possibly obtained using the harmonic balance method [28, 3, 32], linearly perturbed dynamics can be studied to assess stability in the hyperbolic case. Therefore, a Floquet-type theory on LTP-FODEs needs to be developed and the existence of Floquet exponents needs to be studied. For application of this theory to numerical stability analysis, methods like the Hill method should be extended to the fractional-order case.
@@
The main merit of this paper is the insight that exponentially decaying Floquet solutions (i.e. the ones related to asymptotically stable equilibria or stable manifolds of saddle equilibria) do not exist for LTP-FODEs, but exponentially growing Floquet solutions may be observed and assessed via a Hill-type approach. Parts of the research of this work were presented in a preliminary form in reference [15], namely the fractional Hill problem was presented as well as results on linear time-invariant FODEs. The novel contributions of the present work are the formulation of the system class in Section 3, the analysis of decaying solutions in Section 4 as well as the proof of Theorem 5.1 and the following Propositions.
@@ 2.1 Fractional Calculus
We briefly recall some concepts of fractional calculus and refer the reader to the books [10, 24] for further details. Several definitions of fractional derivatives exist, each suited to different mathematical and physical contexts. Given a function $f\in L^{1}[a,b]$, the Riemann–Liouville (RL) fractional integral of order $\alpha>0$ is defined as
@@
$\left({}^{\mathrm{RL}}\operatorname{D}_{a}^{\alpha}f\right)(t)=\left(\operatorname{D}^{1}{I_{a}^{1-\alpha}}f\right)(t).$ (2)
@@
To overcome this limitation of the RL derivative, the Caputo derivative is often used in applied settings [18, 28, 3]. For $\alpha\in(0,1)$ and a function $f\in AC[a,b]$, it is defined as
@@
A different formulation (see [10] and the references therein) of the Caputo derivative reads
@@
for the case of $\alpha\in(0,1)$. If $f\in AC[a,b]$, then the two formulations coincide $\operatorname{D}_{\ast a}^{\alpha}f={}^{\mathrm{C}}\operatorname{D}_{a}^{\alpha}f$ almost everywhere [10].
@@ 2.2 The Hill Method for First-Order Systems
Hill’s method is a frequency-based approach to assess the stability of periodic solutions [38]. It has been applied to first order ordinary differential equations as well as delay differential equations [22]. Here, we summarize the approach for ordinary differential equations, which we develop for FODEs in Section 5. For time-periodic first-order systems
@@
Herein, $\tilde{J}_{k}$ are the Fourier coefficients of $\tilde{J}(t)=\sum_{k=-\infty}^{+\infty}\tilde{J}_{k}\exp(ik\omega t)$, $\omega=2\pi/T$ and $I$ is the identity matrix of appropriate dimension. After truncation to finite dimension $N$, one may either compute all eigenvalues of the truncated Hill-matrix $\mathbf{H}_{N}$ and employ a sorting technique to extract Floquet exponents or perform the recently introduced Koopman-Hill projection [2], avoiding sorting and directly approximating Floquet-multipliers.
@@ 3 Liouville-Weyl Differential Equations: A Class of Systems with Periodic Solutions
$X(-\infty,t_{0}]:=$ $\Bigl\{x_{0}\in{AC_{\mathrm{loc}}(-\infty,t_{0}]}\cap L^{\infty}(-\infty,t_{0}]~|~\exists\,\eta>0:x_{0}^{\prime}\in L^{\infty}[t_{0}-\eta,t_{0}]\Bigr\}$ (13)
@@
${}^{\mathrm{C}}\operatorname{D}_{t_{0}}^{\alpha}x(t)$ $=f(t,x(t))-\mathcal{F}x_{0}(t)=:f_{0}(t,x(t)),\quad t\geq t_{0},\quad x(t_{0})=x_{0}(t_{0}).$ (18)
Typically, the $\operatorname{D}_{\ast 0}^{\alpha}$-type differential equations are considered instead of $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}-$type formulations, since the latter shows problems on the space $C[0,h]$, but the two formulations coincide on $AC[0,h]$, see [36, Section 2.3]. We formalize this for our case in the Proposition 3.5 below. Consider the following example of $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}-$type problems on $C[0,h]$.
@@
Let $x:[0,1]\to\mathbb{R}$ be the Cantor ternary function [11, pp. 96ff.]. This function is continuous, monotonically increasing, differentiable almost everywhere, and satisfies $x^{\prime}(t)=0$ a. e. Thus, $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}x(t)=0$ for $t\in[0,1]$ and $x$ is a strong solution of the initial value problem $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}x(t)=0$, $x(0)=0$. By a strong solution we mean a continuous function satisfying the differential equation and initial condition.
@@
$\operatorname{D}_{\ast t_{0}}^{\alpha}x(t)$ $=f(t,x(t))-\mathcal{F}x_{0}(t)=:f_{0}(t,x(t)),\quad t\geq t_{0},\quad x(t_{0})=x_{0}(t_{0}).$ (19)
Since $f$ is continuous in its first variable and locally Lipschitz in its second variable, the same properties hold for $f_{0}$, because Lemma 3.3 guarantees that the forcing term $\mathcal{F}x_{0}$ is continuous. We thus have existence of strong continuous solutions in the sense of Theorem 6.5 of [10]. These solutions may or may not coincide with solutions of the $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}-$type problem (18). Of course, since the two operators $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}$ and $\operatorname{D}_{\ast 0}^{\alpha}$ coincide almost everywhere in the case of absolutely continuous functions, we have the following statement.
@@
Consider the Weierstrass function [30]
@@
which is continuous but nowhere differentiable on $[0,1]$. This function $W$ is Hölder continuous with exponent $\beta$ for all $\beta\in(0,1)$ and $W(0)=0$. Following the results of [30], $W$ thus has a continuous Caputo fractional derivative $\operatorname{D}_{\ast 0}^{\alpha}W$ of order $\alpha$ for all $\alpha\in(0,1)$. Then, the function $W$ can be seen as a solution of the initial-value problem
@@
The example (i) illustrates that exponentially decaying initial functions are not feasible, as these grow unboundedly in negative time. We will also observe in the next section that decaying solutions of linear versions of (12a) possess an algebraic decay instead of an exponential decay. Exponential growth in case (ii) is indeed possible and will be observed in the two following sections. In example (iii), the initialization function $x_{0}$ is a typical trajectory generated by a linear fractional-order Caputo differential equation with finite lower bound. Although its first derivative is undefined at time t=1t=-1, its forcing term is well defined. Example (iv) is an example showing that a singularity of $x_{0}^{\prime}$ at $t_{0}$ can lead to nonexistence of $\mathcal{F}x_{0}$. This particular example of $x_{0}$ is bounded and uniformly continuous, showing that the space of initial functions of (12a) needs to be chosen differently than in [18]. Example (v) allows for an explicit computation of the forcing term. As this work is motivated by the study of periodic solutions, example (vi) is a periodic initial function. In fact, the Liouville-Weyl differential equation (12a) can possibly posess a periodic solution, in contrast to fractional differential equations with finite lower bound [34]. Indeed, a fractional-order Caputo derivative with finite lower bound of a periodic function is not periodic [1]. A minimal example of a system with a periodic solution is given in the following.
@@
We intend to study stability properties of linear cases of the system (12a) in the subsequent sections. We give the stability definition here, which is equivalent to the one of [18] and the classical literature on functional differential equations [16].